OpenStudy (help_needed_cudi):

7. Is (-7, 4) a solution to the equation y = x2 – 3? (1 point) Yes No

OpenStudy (31356):

Go on your graphing calculator :D

OpenStudy (texaschic101):

just sub in -7 for x and 4 for y y = x^2 - 3 4 = -7^2 - 3 think you can finish ?

OpenStudy (31356):

I think your equation is y=x^2.... right?

OpenStudy (texaschic101):

y = x^2 - 3

OpenStudy (31356):

Okay :D

OpenStudy (31356):

As a result, It's (-7, 46) which means it's not a solution :D